# https://www.intelligence.tuc.gr/gogs/

### v2

@@ 6 Open Challenges
As we have stressed earlier in this deliverable, the main challenges for applying multiagent systems that are able to learn and adapt in real applications are the following:
@@
  • Uncertainty and heterogeneity with respect to the environment, the agent beliefs, and the agent internal structures. Sophisticated solutions should account for probabilistic formulations that are able to capture the inherent uncertainty; and also allow for fitting heterogeneous models, as uniformity is rarely found in practice.
@@
  • Game-theoretic analysis and tractability. The convergence of a successful learning process must be guaranteed and the solutions should have desired properties considering agent strategies and incentives. The setting of respective bounds is possible via GT analysis [yang2020overview].
@@
For PTM, other, novel applications, different than the conventional formulations for document analysis, can be proven promising and valuable. Overall, the combination of GNNs, DRL, and PTM, being applied to tackle requirements that are complementary to each other, is a promising direction to consider. As the survey of [Fathinezhad23] also suggests, combinations of GNN and DRL lead to realistic and effective solutions for real-world problems.
Finally, with respect to the domain of GT, there is the need to redefine equilibrium concepts in a way that enables practical strategic decision-making in complex environments, surpassing the limitations of traditional GT assumptions. By integrating heterogeneous agent beliefs into the computation process, novel equilibrium concepts with bounded rationality and flexible performance guarantees may arise. For this, our ultimate aim is to blur the boundary between cooperative and non-cooperative game theory, and to offer novel solutions applicable across various domains.
@@ 7 Conclusions
Under this prism, we reviewed methods and approaches in the domains of graph neural networks, deep reinforcement learning, and probabilistic topic modeling. We presented the most known GNN leargning algorithms variations, and referred to past work that has applied them to multiagent strategic settings. We distinguished among approaches that consider spatial, spectral, and sequential aspects, as well as those that require assumptions related to common priors or agent self interest. With respect to DRL, we introduced the required background for tabular and single-agent DRL and delved into multiagent-enabled algorithms. We analytically described the related work and highlight particular approach characteristics such as decentralization, training philosophy, and inter-agent communication or self interest requirements. Further, we discussed the theoretical basis of the well-known solution concepts such as the Nash equilibrium and the notion of correlated equilibria, as well as desired game theoretic properties such as fairness and stability. The field of PTM is also interesting due to its capability to fit unknown underlying distributions, such as the underlying structures governing agent beliefs and types. We presented existing work that utilizes PTM in domains other than document classification, for which this was originally designed for. Finally, we outlined certain open challenges that exist in this highly challenging research domain.
@@ 7.0.1 Acknowledgements
The research described in this paper was carried out within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union - NextGenera- tionEU (Implementation Body: HFRI. Project name: DEEP- REBAYES. HFRI Project Number 15430).

### v3

@@ 6 Open Challenges
As stressed early on in this paper, in our view the main challenges for applying multiagent systems that are able to learn and adapt in real-world applications are the following:
@@
  • Uncertainty and heterogeneity with respect to the environment, agent beliefs, agent capabilities, and agent internal structures. Sophisticated solutions should account for probabilistic formulations that are able to capture the inherent uncertainty; and also allow for fitting heterogeneous models, as uniformity is rarely found in practice.
    • We note that a recent framework put forward by Bakopoulos and Chalkiadakis [9] allows for the principled handling of uncertainty to increase exploration efficiency. This adaptive exploration framework prevents the agent from over-exploring previously explored parts of the environment, instead encouraging exploration of states with higher uncertainty. To create an instance of this framework, a user should select (among others) a definition of uncertainty. In addition, all previous adaptive exploration methods can be viewed as special cases of this framework, when initialized with the appropriate definition of uncertainty.
    • In addition, as already mentioned earlier, we note that a AAAI-2025 publication by Chrysomallis and Chalkiadakis [28] introduces a novel, heterogeneous actions variant (HA-DIIQN) of the DIIQN algorithm, enabling implicit imitation between agents with distinct action spaces, and thus broadening the applicability of deep implicit imitation RL to real-world scenarios with diverse agent characteristics.
@@
  • Game-theoretic analysis and tractability. The convergence of a successful learning process must be guaranteed and the solutions should have desired properties considering agent strategies and incentives. The setting of respective bounds is possible via GT analysis [155].
@@
For PTM, other, novel applications, different than the conventional formulations for document analysis, can be proven promising and valuable. Overall, the combination of GNNs, DRL, and PTM, being applied to tackle requirements that are complementary to each other, is a promising direction to consider. As the survey of [39] also suggests, combinations of GNN and DRL lead to realistic and effective solutions for real-world problems.
Finally, with respect to the domain of GT, there is the need to (re-)define equilibrium concepts in a way that enables practical strategic decision-making in complex environments, surpassing the limitations of traditional GT assumptions. By integrating heterogeneous agent beliefs into the computation process, novel equilibrium concepts with bounded rationality and flexible performance guarantees may arise—with an ultimate goal to blur the boundary between cooperative and non-cooperative game theory, and to offer novel solutions applicable across various domains.
@@ 7 Conclusions
Under this prism, we reviewed methods and approaches in the domains of graph neural networks, deep reinforcement learning, and probabilistic topic modeling. We presented the most known GNN leargning algorithms variations, and referred to past work that has applied them to multiagent strategic settings. We distinguished among approaches that consider spatial, spectral, and sequential aspects, as well as those that require assumptions related to common priors or agent self interest. With respect to DRL, we introduced the required background for tabular and single-agent DRL and delved into multiagent-enabled algorithms. We provided an overview related work and highlighted particular approach characteristics such as decentralization, training philosophy, and inter-agent communication or self interest requirements. Further, we discussed the theoretical basis of the well-known solution concepts such as the Nash equilibrium and the notion of correlated equilibria, as well as desired game theoretic properties such as fairness and stability. The field of PTM is also interesting due to its capability to fit unknown underlying distributions, such as the underlying structures governing agent beliefs and types. We presented existing work that utilizes PTM in domains other than document classification, for which this was originally designed for. Finally, we outlined certain open challenges that exist in this highly challenging research domain.
@@ 7.0.1 Acknowledgements
The research described in this paper was carried out within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union - NextGenerationEU (Implementation Body: HFRI. Project name: DEEP- REBAYES. HFRI Project Number 15430). We also thank Stefano Albrecht for useful comments on an earlier version of this paper.